OpenStudy (anonymous):

integration of (squareroot of cos 2x )/sinx wrt dx?

OpenStudy (anonymous):

\[\int\limits_{?}^{?} (\sqrt{\cos 2x})\div (\sin x)*dx\]

OpenStudy (dan815):